What is the value of x in the given system of equations: y = 3x + 1 and y = 5x - 3
Show work.
A)1
B)7
C)2
D)-3

Answer by mananth(16946)

y = 3x + 1 and y = 5x - 3
substitute y in the second equation

3x+1=5x-3
3x-5x=-3-1
-2x=-4
/-2
x=2